OpenStudy (anonymous):

2/3 k + 6 = −12 for k

OpenStudy (anonymous):

i got -3 just checking if im right ...

OpenStudy (anonymous):

Order of operations... Subtract 6 from both sides; then divide by (2/3), which is the same thing as multiplying by 3/2

jhonyy9 (jhonyy9):

2k+18=-36 2k=-54 k=-27

OpenStudy (anonymous):

jhonyy9 did it another way, he multiplied by 3 first to everything

OpenStudy (anonymous):

Same answer, -27

OpenStudy (anonymous):

so is he right ?

OpenStudy (anonymous):

We both are. There's often more than one way to go about a math problem you know ;-)

OpenStudy (anonymous):

\[\frac{2}{3}k+6=-12\] \[\frac{2}{3}k=-18\] \[k=-18*\frac{3}{2}=-27\]